How to find the value of cos 37°?with full procedure pelase let me know ??????????

Grade:11

5 Answers

Harsh Patodia IIT Roorkee
7 years ago
U have to look for trigonometric table to find the values. We can only derive for certain values.
Shreyas Patil
7 years ago
You can get the values of all trigonometry functions in log book. Purchase it from any local book store.
Hope it helps.
Thank You.
Manas Shukla
7 years ago
U need to find from table although
37 degrees is a comonly occuring in questions so just remember the following
37 degree is an angle of a right angle triangle whose sides are 3units. 4 units and 5 units whose base is 4, hypotenuse 5 and perpendicular is 3.
Thus
sin 37 = 3/5
cos 37 = 4/5 and
tan 37 = 3/4
I hope that helps.
 
Piyush Kumar Behera
7 years ago
You should not be concerned about such things,take it as granted that these are the angles in a right triangle having side of the form 3n,4n,5n,n\inN.
Harshit Panwar
6 years ago
Refer log book.........
 
But my teacher taught me by taking an example of a triangle of side 5 units as hypotenuse and 4 &3 as other sides..